OpenStudy (anonymous):

let r = cos(theta) + sin(theta) a. show that the equation in polar coordinates is a circle by writing the corresponding rectangular form. b. provide the center and radius anyone have any idea how to do this problem?

OpenStudy (anonymous):

r = cos(theta) + sin(theta) multipy both sides by r and use r^2=x^2+y^2 and rcos(theta)=x and rsin(theta)=y thus we have r ^2= rcos(theta) + rsin(theta) or x^2+y^2=x+y or x^2+y^2-x-y=0 or x^2-2(1/2)x +1/4+y^2-2(1/2)y +1/4 =1/4+1/4 or (x-1/2)^2 +(y-1/2)^2= 1/2 circle with centre at(1/2,1/2) and radius=

OpenStudy (anonymous):

r=1/sqrt(2)

OpenStudy (anonymous):

thank you so much!